# Instructions for *ACL Proceedings

### v1

@@ Abstract
Counter-speech (CS) is a key strategy for mitigating online Hate Speech (HS), yet defining the criteria to assess its effectiveness remains an open challenge. We propose a novel computational framework for CS effectiveness classification, grounded in social science concepts. Our framework defines six core dimensions – Clarity, Evidence, Emotional Appeal, Rebuttal, Audience Adaptation, and Fairness – which we use to annotate 4,214 CS instances from two benchmark datasets, resulting in a novel linguistic resource released to the community. In addition, we propose two classification strategies, multi-task and dependency-based, achieving strong results (0.94 and 0.96 average F1 respectively on both expert- and user-written CS), outperforming standard baselines, and revealing strong interdependence among dimensions.
Effectiveness of Counter-Speech against Abusive Content: A Multidimensional Annotation and Classification Study
Greta Damo1, Elena Cabrio1 Serena Villata1 1Université Côte d’Azur, CNRS, Inria, I3S, France Correspondence: greta.damo@univ-cotedazur.fr
@@
1 Introduction
The rise of social media has intensified concerns over online Hate Speech (HS). Counter-speech (CS), non-aggressive responses that challenge hate using facts or alternative views, has emerged as a promising mitigation strategy Schieb and Preuss (2016); Benesch (2014). CS has been studied through various lenses: some see it as direct responses to hate Mathew et al. (2019); Ashida and Komachi (2022), while others argue it should condemn hate and offer victims support and alternative views Vidgen et al. (2021); Hangartner et al. (2021); Vidgen et al. (2020); He et al. (2021). CS is also valued for its potential to shift beliefs and reduce future hate Qian et al. (2019a); Rieger et al. (2018). CS can be written by experts, online users, or chatbots (based on Large Language Models – LLMs). Expert-written CS typically presents structured reasoning and appeals to shared human values, whereas user and LLM responses tend to be more generic and emotional Mun et al. (2023); Tekiroğlu et al. (2022, 2020).
This study evaluates expert- and user-written CS effectiveness using two benchmark datasets Albanyan and Blanco (2022); Chung et al. (2019). Figure 1 shows examples of expert and user CS.
@@ 1 Introduction
Current evaluations on CS effectiveness rely on surface-level metrics – coherence, relevance, or non-offensiveness – assuming well-formed CS is inherently effective Chung et al. (2021); Zhu and Bhat (2021); Baheti et al. (2021); Ashida and Komachi (2022); Mun et al. (2023). But concentrating on CS surface form only is not enough to achieve shifting attitudes or reducing harm.
@@
To address this gap, our contribution is twofold: (1) a new multi-dimensional framework for automatic CS effectiveness classification across six dimensions – Clarity, Evidence, Emotional Appeal, Rebuttal, Audience Adaptation, and Fairness – from linguistics and argumentation theories, and two classification strategies, multi-task and dependency-based modeling, outperforming standard baselines, achieving average F1 scores of 0.94 and 0.96 respectively, when combining user-written and expert-written CS data. (2) a novel human annotated resource combining two expert- Chung et al. (2019) and user-written Albanyan and Blanco (2022) CS datasets, enriched with the annotation layer with the six dimensions [^1] Datasets, guidelines, and code are available at this link..
2 Related Work
CS, introduced as a mitigation strategy by Susan et al. (2016), has not yet been extensively explored from a detection standpoint. Among the studies, Poudhar et al. (2024) find that a one-shot prompted LLM achieves promising accuracy in classifying manually labeled CS strategies. Mathew et al. (2018) analyze HS/CS tweet pairs, using a boosting algorithm with TF-IDF and lexical features, which achieves 0.77 F1. Mathew et al. (2019) also analyze CS and neutral comments on YouTube, finding that TF-IDF vectors combined with logistic regression achieve 0.73 F1. Ziems et al. (2020) analyze tweets during the COVID-19 crisis, achieving 0.49 F1 for CS. Garland et al. (2020) apply an ensemble learning algorithm to classify HS and CS tweets, achieving F1 scores between 0.76 and 0.97. A key gap in the literature is the lack of standardized evaluation methods for assessing CS. Several studies propose evaluation criteria, but no unified framework emerged. Zheng et al. (2023) assess CS effectiveness through human judgments, focusing on comfort for targets and bystander empathy. Chung et al. (2021) and Zhu and Bhat (2021) similarly rely on human ratings based on qualities such as informativeness, coherence, and stance. Ashida and Komachi (2022) evaluate CS suitability using offensiveness and stance, while Jones et al. (2024) propose an LLM-based evaluation framework, grounded in NGO guidelines, which shows strong alignment with human judgments. Yu et al. (2024) introduce a conversation-level incivility metric, arguing that CS should be judged by its downstream impact on discourse. We advance beyond previous studies by introducing a human-annotated resource that captures fine-grained dimensions of structural and rhetorical message quality, and by developing classification methods that leverage this framework.
3 Effectiveness Metrics
We propose a framework to evaluate CS effectiveness across six dimensions: Clarity, Evidence, Emotional Appeal, Rebuttal, Audience Adaptation, and Fairness. Clarity ensures structural coherence and logical flow, allowing to follow texts effectively Persing and Ng (2013); Wachsmuth et al. (2016, 2017); Stapleton and Wu (2015). Evidence strengthens persuasiveness through relevant support such as statistics or expert testimony Wachsmuth et al. (2017); Rinott et al. (2015); Rahimi et al. (2014); Reinard (1988). Emotional Appeal, a rhetorical strategy, shows that empathy and emotion can enhance impact and share Wachsmuth et al. (2017); Aristotle (2007); Govier (2010); Hangartner et al. (2021); Chen et al. (2022); Marone (2015). Rebuttal reflects the ability to address and counter opposing views, improving credibility and persuasion Wachsmuth et al. (2017); Aristotle (2007); Toulmin (1958); Damer (2009); Granger et al. (2009); O’Keefe (1999); Onoda et al. (2015). Audience Adaptation involves tailoring language to the audience’s style and knowledge, improving relatability Pickering and Garrod (2004); Danescu-Niculescu-Mizil et al. (2011); Wang et al. (2014). Fairness promotes respectful tone and avoids dehumanization, supporting ethical discourse Cialdini (1993); Schreier et al. (1995).
@@ 3 Effectiveness Metrics
Table 1: Annotation guidelines for CS effectiveness.
@@
Emotional Appeal and Audience Adaptation are binary (0–1) variables, where 1 indicates presence and 0 absence. Similar to prior work simplifying subjective dimensions into binary formats Habernal and Gurevych (2016); Wachsmuth et al. (2017); Oraby et al. (2017); Koszowy et al. (2024), we adopt this approach to reduce ambiguity. Clarity, Evidence, Rebuttal, and Fairness are categorical, using a 1–3 Likert scale, with 3 as the best score. These dimensions rely on more objective indicators (for example, the number of evidence pieces) that capture weak, moderate, and strong instances. Table 1 reports the definitions for the annotation of the effectiveness dimensions.
4 Datasets Description
@@
We extend the annotation on two existing datasets: CONAN Chung et al. (2019) and the Twitter Dataset Albanyan and Blanco (2022). We select CONAN as it is the first expert-curated dataset of HS/CS pairs, widely used as a benchmark. It is a multilingual (English, French, Italian) dataset centered on Islamophobia, containing 4,078 expert-annotated HS/CS pairs. Through translation and paraphrasing, it is expanded to 14,988 pairs. For our experiments, we retain only the English, non-augmented instances, resulting in 3,847 pairs. The Twitter Dataset [^2] Only tweet IDs are publicly available due to privacy policies. We obtained the full data upon request from the authors. contains 5,652 hateful tweets and replies from social media (Twitter/X), capturing the brevity and style typical of online discourse Bonaldi et al. (2024). We focus on the subset labeled as CS, where a clear target is identifiable, yielding 367 HS/CS pairs. Together, these datasets form a combined benchmark of 4,214 HS/CS pairs.
@@
Annotation & IAA. We define annotation guidelines to label CS effectiveness dimensions. To consolidate these guidelines, a pilot study was conducted on 50 HS/CS pairs from each dataset, involving three annotators  [^3] Two female, one other; age group: 21–30; education level: PhD students. with background in computer science. Through different reconciliation phases, the guidelines were improved iteratively. After an initial round, annotators reviewed disagreements to resolve confusion or misinterpretation. Based on their feedback, the guidelines were updated, and a second and final annotation round was conducted. To measure inter-annotator agreement (IAA), Fleiss’s $\kappa$ was used for binary dimensions, Krippendorff’s $\alpha$ for categorical ones, and Percent Agreement for Audience Adaptation due to near-perfect consensus (as other metrics underestimate high non-random agreement). IAA results improved significantly in the second round, after which the final dataset was labeled by one of the annotators. Table 2 shows IAA scores, indicating strong agreement across all dimensions (see Appendix A for pairwise scores). We also show average effectiveness scores, demonstrating that expert-written CS is more effective (avg. 1.62) than user-written CS. CONAN scores higher in all dimensions except emotional appeal, suggesting online users rely more on emotions than evidential reasoning or facts.
@@
Table 2: IAA from two annotation rounds (Fleiss’ $\kappa$ for binary variables, and Krippendorff’s $\alpha$ for categorical ones) and average effectiveness scores.
5 Experimental Setting
We use BERT-based classifiers to predict CS effectiveness scores across the six dimensions, relying on bert-base-uncased Devlin et al. (2018) trained on the combined dataset of 4,214 HS/CS pairs, and cross-dataset evaluations. We use BERT as the sole model to focus on improving it through our framework. By not comparing across different architectures, we isolate the effect of our multi-task and dependency modeling, ensuring that observed gains result from our contributions. We assess the following configurations (Appendix B for details):
Bert_CS: CS text embeddings are used as input. Binary dimensions are trained using Binary Cross-Entropy (BCE) loss function, and categorical ones with Cross-Entropy (CE). Bert_CS_HS: identical to Bert_CS, but the input is the concatenation of HS and CS texts embeddings for more context. Multi-task_divided trains binary dimensions jointly (BCE for Emotional Appeal, Focal Loss for Audience Adaptation) and categorical ones separately (CE), assuming partial correlation. Multi-task_united combines all dimensions in a single model using the sum of their respective loss functions, assuming stronger interdependence. Dependency_matrix_3e introduces a learnable matrix to capture pairwise relationships between dimensions, and is trained for 3 epochs, while Dependency_matrix_6e for 6 epochs to assess the effect of extended training.
6 Results
Model Emo. Aud. Clarity Evid. Rebut. Fair. AVG
Train: Combined Dataset
bert_cs 0.67±0.09 0.99±0.00 0.94±0.01 0.96±0.01 0.96±0.01 0.95±0.01 0.91±0.02
bert_cs_hs 0.62±0.12 0.99±0.00 0.94±0.02 0.96±0.01 0.95±0.01 0.94±0.01 0.90±0.02
multi-task_d 0.92±0.02 0.99±0.00 0.92±0.01 0.93±0.01 0.94±0.03 0.92±0.03 0.94±0.01
multi-task_u 0.84±0.04 0.99±0.00 0.93±0.02 0.92±0.02 0.93±0.03 0.93±0.02 0.92±0.01
dep._m._3e 0.80±0.10 0.99±0.00 0.92±0.01 0.90±0.05 0.91±0.03 0.93±0.03 0.91±0.02
dep._m._6e 0.92±0.04 0.99±0.00 0.95±0.01 0.97±0.02 0.96±0.01 0.95±0.02 0.96*±0.01
Train: CONAN $\rightarrow$ Test: Twitter
bert_cs 0.30±0.01 1.00±0.00 0.41±0.04 0.86±0.09 0.77±0.02 0.54±0.21 0.65±0.05
bert_cs_hs 0.30±0.04 1.00±0.00 0.33±0.03 0.84±0.07 0.73±0.05 0.41±0.07 0.60±0.02
multi-task_d 0.49±0.07 1.00±0.00 0.42±0.02 0.91±0.04 0.80±0.02 0.32±0.03 0.65±0.02
multi-task_u 0.52±0.06 1.00±0.00 0.42±0.03 0.93±0.04 0.77±0.02 0.35±0.05 0.66±0.01
dep._m._3e 0.54±0.02 1.00±0.00 0.41±0.02 0.92±0.02 0.77±0.02 0.34±0.05 0.66±0.01
dep._m._6e 0.50±0.05 1.00±0.00 0.42±0.01 0.92±0.03 0.77±0.02 0.38±0.01 0.66*±0.01
Train: Twitter $\rightarrow$ Test: CONAN
bert_cs 0.14±0.00 0.99±0.00 0.74±0.05 0.45±0.06 0.55±0.11 0.92±0.01 0.63±0.03
bert_cs_hs 0.14±0.01 0.99±0.00 0.76±0.06 0.43±0.02 0.55±0.14 0.92±0.02 0.63±0.03
multi-task_d 0.55±0.17 0.99±0.00 0.51±0.24 0.44±0.01 0.60±0.05 0.89±0.13 0.66*±0.04
multi-task_u 0.47±0.23 0.99±0.00 0.38±0.27 0.43±0.02 0.56±0.07 0.72±0.39 0.59±0.09
dep._m._3e 0.40±0.24 0.99±0.00 0.48±0.22 0.43±0.03 0.53±0.09 0.70±0.13 0.59±0.08
dep._m._6e 0.40±0.21 0.99±0.00 0.75±0.04 0.53±0.07 0.47±0.21 0.78±0.14 0.65±0.07
Table 3: F1 scores (± std) for all models with CS embeddings. Statistically significant results are marked with *.

### v2

# Effectiveness of Counter-Speech against Abusive Content: A Multidimensional Annotation and Classification Study

@@ Abstract
Counter-speech (CS) is a key strategy for mitigating online Hate Speech (HS), yet defining the criteria to assess its effectiveness remains an open challenge. We propose a novel computational framework for CS effectiveness classification, grounded in linguistics, communication and argumentation concepts. Our framework defines six core dimensions – Clarity, Evidence, Emotional Appeal, Rebuttal, Audience Adaptation, and Fairness – which we use to annotate 4,214 CS instances from two benchmark datasets, resulting in a novel linguistic resource released to the community. In addition, we propose two classification strategies, multi-task and dependency-based, achieving strong results (0.94 and 0.96 average F1 respectively on both expert- and user-written CS), outperforming standard baselines, and revealing strong interdependence among dimensions.
@@
I Introduction
The rise of social media has intensified concerns over online Hate Speech (HS). Counter-speech (CS) has emerged as an effective strategy for mitigating harmful online discourse. It involves non-aggressive responses, that use credible evidence, factual arguments, and alternative perspectives to challenge hate speech [1, 2]. Counter-speech has been studied and approached through various lenses: some see it as direct responses to hate speech [3, 4], while others argue it should condemn hate, challenge and offer alternative viewpoints, and offer support to the victim [5, 6, 7, 8]. CS is also valued for its potential to shift beliefs and reduce future hate [9, 10]. CS can be written by experts, online users, or chatbots (based on Large Language Models – LLMs). Expert-written CS typically presents structured reasoning and appeals to shared human values, whereas user- and LLM-generated responses, in particular, tend to be more generic and emotional, and often lacks depth and relies on surface-level rebuttals [11, 12, 13].
This study evaluates expert- and user-written CS effectiveness using two benchmark datasets [14, 15].This allows us to compare how the origin of a counter-speech message influences its effectiveness. Figure 1 shows an example of a hateful message with user- and expert-written CS responses.
@@ I Introduction
Current evaluations of counter-speech effectiveness rely on surface-level metrics – coherence, relevance, or non-offensiveness – assuming well-formed CS is inherently effective [16, 17, 18, 4, 11]. But concentrating on CS surface form only is not enough to achieve shifting attitudes or reducing harm. Assessing CS effectiveness is crucial, as it has the potential for significant societal impact, by reducing online hostility and potentially curbing offline violence and discrimination.
@@
To address this gap, this study introduces a novel structured framework grounded in linguistics, communication, rhetoric, and argumentation theories, helping to evaluate counter-speech effectiveness. Our contribution is twofold: (1) a new multi-dimensional framework for automatic counter-speech effectiveness classification across six key dimensions – Clarity, Evidence, Emotional Appeal, Rebuttal, Audience Adaptation, and Fairness – from linguistics, communication, rhetoric and argumentation theories, and two classification strategies, multi-task and dependency-based modeling, that outperform standard baselines, achieving average F1 scores of 0.94 and 0.96 respectively, when combining user-written and expert-written CS data. (2) a novel human annotated resource comprising two expert- [15] and user-written [14] counter-speech datasets, enriched with the annotation layer with the six dimensions of our effectiveness framework [^1] Datasets, guidelines, and code are available at this link..
The remainder of the paper is structured as follows. Section II reviews prior studies on counter-speech and its evaluation. Section III defines the six dimensions used to assess CS effectiveness, tailoring them to existing linguistics and argumentation theories. Section IV details the datasets used for the annotation and the classification task, and the annotation process together with the Inter-Annotator-Agreement results. Section V outlines the models used, and the experimental setup. Section VI presents the results of our classification task, showing comparisons between the baselines and our proposed models. Section VII identifies and discusses challenges in specific dimensions and limitations. Finally, Section VIII summarizes key findings and future directions.
II Related Work
Research on counter-speech has gained increasing attention as a strategy for combating online hate speech. However, most studies have focused on hate speech detection and counter-speech generation, while the systematic evaluation of CS effectiveness remains under-explored. In this section, we review key contributions on counter-speech analysis, emphasizing the need for a comprehensive CS evaluation framework.
II-A Hate Speech & Counter Speech Detection
HS detection is a well-studied NLP task, initially tackled as binary classification on English datasets [19, 20], later extended to fine-grained labels for discrimination and microaggressions [21, 22, 23, 24, 25, 26]. Early approaches relied on traditional Machine Learning [27, 28], while recent methods leverage transformers like BERT for better contextual modeling [29, 30, 31, 32].
On the other hand, counter-speech, introduced as a mitigation strategy by [33], has not yet been extensively explored from a detection standpoint. Among the studies, [34] propose an early HS/CS detection pipeline, by applying an ensemble learning algorithm to classify HS and CS tweets, achieving F1 scores between 0.76 and 0.97, emphasizing counter-speech complexity due to subjectivity and definitional ambiguity.
II-B Counter Speech Classification & Evaluation
Concerning counter-speech classification, [35] find that a one-shot prompted LLM achieves promising accuracy in classifying manually labeled CS strategies. [36] analyze HS/CS tweet pairs, using a boosting algorithm with TF-IDF and lexical features, which achieves 0.77 F1. [3] also analyze CS and neutral comments on YouTube, finding that TF-IDF vectors combined with logistic regression achieve 0.73 F1. [37] conducts a qualitative analysis of counter-speech examples, while [38] analyze tweets during the COVID-19 crisis, achieving an F1 score of 0.49 for CS.
A key gap in the literature is the lack of standardized evaluation methods for assessing human- or machine-generated counter-speech. Several studies propose evaluation criteria, but no unified framework emerged. [39] assess CS effectiveness through human judgments, focusing on comfort for targets and bystander empathy. [16] and [17] similarly rely on human ratings based on qualities such as informativeness, coherence, and stance. [4] evaluate CS suitability using offensiveness and stance, while [40] propose an LLM-based evaluation framework, grounded in NGO guidelines, which shows strong alignment with human judgments. [41] introduce a conversation-level incivility metric, arguing that CS should be judged by its downstream impact on discourse. Finally, [42] focus on evaluating the human-likeness of AI-generated counter-speech.
While prior work shows growing interest in CS evaluation, we advance beyond previous studies by introducing a human-annotated resource based on our framework that captures fine-grained dimensions of structural coherence, content strength, rhetorical impact, and linguistic strategies, to assess its effectiveness. We also develop classification methods that leverage this framework.
III Effectiveness Metrics
We propose a framework to evaluate counter-speech effectiveness across six dimensions: Clarity, Evidence, Emotional Appeal, Rebuttal, Audience Adaptation, and Fairness. These are grounded in communication theory.
  • Clarity ensures structural coherence and logical flow. Effective counter-speech should be clear and logically structured, so that it is easier for the audience to follow the reasoning and understand the main points [43, 44, 45, 46].
  • Evidence ensures that effective counter-speech is supported by relevant evidence and examples that support the claims and make it more compelling to the audience. They should be specific to the topics addressed in the hateful message. Studies indicate that presenting multiple pieces of evidence in an argument, such as statistics or expert testimony, enhances persuasiveness and effectiveness [45, 47, 48, 49].
  • Emotional Appeal is a rhetorical strategy, used to evoke empathy, or other emotions in the audience, which can help strengthen the counter-speech impact. The importance of emotions and empathy in designing effective counter-speech has been investigated by numerous studies [45, 50, 51, 6, 52, 53].
  • Rebuttal shows how using arguments to anticipate and address potential counterarguments, objections, and opposing views improves credibility and strengthens persuasion [45, 50, 54, 55, 56, 57, 58].
  • Audience Adaptation involves tailoring the language of the counter-speech to a specific audience, taking into account their level of linguistic ability and knowledge, thus improving understanding and relatability. From a linguistic point of view, a counter-speech is more effective and understandable if the language employed is similar to the one spoken by the audience [59, 60, 61].
  • Fairness promotes the use of appropriate and respectful language, respecting freedom of expression, without censoring or dehumanizing the opposing viewpoint. Fair language supports ethical discourse and improves effectiveness [62, 63].
Unlike prior frameworks, focused narrowly on toxicity or engagement, our approach enables fine-grained analysis informed by the theoretical perspectives described in argumentation and communication studies. They are divided into binary and categorical metrics.
TABLE I: Annotation guidelines for CS effectiveness.
@@ III Effectiveness Metrics
Binary metrics. Emotional Appeal and Audience Adaptation are binary (0-1), where 1 indicates presence and 0 absence of the dimension. Due to their inherent subjectiveness and context-dependency, a binary classification (presence vs. absence) seems reasonable. Similar to prior work simplifying subjective dimensions into binary formats [64, 45, 65, 66], we adopt this approach to reduce ambiguity and subjectivity.
@@
Categorical metrics. Clarity, Evidence, Rebuttal, and Fairness are categorical dimensions, using a 1–3 Likert scale, with 3 as the best score. Compared to the binary variables, these dimensions rely on more objective indicators (for example, the number of pieces of evidence), therefore, a 1-3 scale allows annotators to capture gradual differences, distinguishing between weak, moderate, and strong instances. Table I reports a summary of the definitions for the annotation of the effectiveness dimensions.
IV Datasets Description
@@
We extend the annotation on two existing datasets: CONAN [15] and the Twitter Dataset [14]. We select CONAN as it is the first expert-curated dataset of HS/CS pairs, widely used as a benchmark. It is a multilingual (English, French, Italian) dataset centered on Islamophobia, containing 4,078 expert-annotated HS/CS pairs. Through translation and paraphrasing, it is expanded to 14,988 pairs. For our experiments, we retain only the English, non-augmented instances, resulting in 3,847 pairs. The Twitter Dataset [^2] Only tweet IDs are publicly available due to privacy policies. We obtained the full data upon request from the authors. is a real-world dataset, containing 5,652 hateful tweets and replies obtained from social media (Twitter/X), capturing the brevity and style typical of online discourse. We focus on the subset labeled as counter-speech, where a clear target is identifiable, yielding 367 HS/CS pairs. Together, these datasets form a combined benchmark of 4,214 HS/CS pairs, used in our experiments.
@@
Annotation & IAA. We define annotation guidelines to label CS effectiveness dimensions. To consolidate these guidelines [^3] Complete guidelines are available at this link, a pilot study was conducted on 50 HS/CS pairs from each dataset, involving three annotators  [^4] Two female, one other; age group: 21–30; education level: PhD students. with background in computer science and linguistics. Through different reconciliation phases, the guidelines were improved iteratively. After an initial round, annotators reviewed disagreements to resolve confusion or misinterpretation. Based on their feedback, the guidelines were updated, and a second and final annotation round was conducted. To measure inter-annotator agreement (IAA), Fleiss’s $\kappa$ was used for binary dimensions, Krippendorff’s $\alpha$ for categorical ones, and Percent Agreement for Audience Adaptation due to near-perfect consensus (as other metrics underestimate high non-random agreement). Table II shows IAA results among the three annotators. The IAA of the second round of annotations improved significantly, compared to the first one, as the annotators had the possibility to discuss and refine their understanding of the guidelines. Overall, the final IAA scores indicate strong agreement across all dimensions, so the remaining part of the dataset was labeled by one of the annotators. Table II also shows average effectiveness scores across both datasets. Expert-written counter-speech obtains an average effectiveness score of 1.62, demonstrating its higher effectiveness compared to user-written counter-speech with an average score of 1.34. CONAN scores higher in all dimensions except emotional appeal, suggesting that online users rely more on emotions than evidential reasoning or facts.
TABLE II: IAA from two annotation rounds (Fleiss’ $\kappa$ for binary variables, and Krippendorff’s $\alpha$ for categorical ones) and average effectiveness scores.
@@
V Experimental Setting
In this section, we describe the model configurations used in our experiments. The goal is to predict CS effectiveness scores across six dimensions. We use BERT classifiers based on bert-base-uncased [67] to predict the effectiveness scores for a given counter-speech response. In all configurations, the max_length is set to 128. The batch_size is 16, and the learning_rate is 2e-5. We run the experiments using five different seeds (0, 1, 2, 3, 42), and we average the results. To perform the experiments, we concatenate both CONAN and Twitter datasets, and then we randomly split it into train, validation, and test subsets, with a percentage of 70-10-20. We also perform cross-validation, by using a single dataset for training and validation, and the other for testing. The experiments were run on one A100 GPU.
We use BERT as the sole model to focus on improving it through our framework. By not comparing across different architectures, we isolate the effect of our proposed models, ensuring that observed gains result from our contributions, rather than differences between model architectures.
We assess the following configurations:
  • Bert_CS: counter-speech text embeddings are used as input. Binary dimensions are trained using the Binary Cross-Entropy (BCE) loss function, and categorical ones with Cross-Entropy Loss (CE).
  • Bert_CS_HS: identical to Bert_CS, but the input is the concatenation of hate speech and counter-speech text embeddings, to add more context.
  • Multi-task_divided. Binary dimensions are trained jointly using BCE (for emotional_appeal) and Focal Loss (to mitigate class imbalance for audience_adaptation), while categorical dimensions are trained separately using CE. Final predictions are obtained by combining all loss components. This setting assumes partial correlation among grouped dimensions.
  • Multi-task_united. All dimensions (binary and categorical) are trained together using their respective loss functions in a single multi-task model. The total loss is computed as the sum of individual losses. This configuration assumes stronger interdependence among all dimensions.
  • Dependency_matrix_3e. All the dimensions are considered together, with their respective loss function (BCE for emotional_appeal, Focal Loss for audience_adaptation, and Cross Entropy Loss for all the multi-label dimensions), which are summed up at the end. Additionally, there is a new parameter, a learnable dependency matrix, to capture pairwise relationships between dimensions. Its rows and columns contain values between 0 and 1, that are weights accounting for the dependency between pairs of dimensions. Its weights are randomly initialized, and their best value is learned during the training phase for 3 epochs.
  • Dependency_matrix_6e has the same configuration of Dependency_matrix_3e, but it is trained for 6 epochs to assess the effect of extended training.
We experiment with two types of input embeddings: (i) concatenated Hate Speech (HS) and Counter Speech (CS) embeddings to provide broader context and (ii) CS embeddings alone. The latter consistently yields slightly better performance. Consequently, we adopt CS-only embeddings for all our proposed configurations (Multi-task_divided, Multi-task_united, Dependency_matrix_3e, and Dependency_matrix_6e).
VI Results
TABLE III: F1 scores (± std) for all models. Statistically significant results are marked with *. AVG: macro-average. Emo.: Emotional Appeal, Aud.: Audience Adaptation, Clarity: Clarity, Evid.: Evidence, Rebut.: Rebuttal, Fair.: Fairness.
Model Train: Combined Dataset Train: Twitter $\rightarrow$ CONAN Train: CONAN $\rightarrow$ Twitter
Emo. Aud. Clarity Evid. Rebut. Fair. AVG Emo. Aud. Clarity Evid. Rebut. Fair. AVG Emo. Aud. Clarity Evid. Rebut. Fair. AVG
[…]